A radio station can be heard 100 miles in all dirctions during evening hours. How many square miles are in the stations broadcast area? use 3.14 as the approximiate value of 3.14
Answer by checkley79(3341)

AREA=PI*R^2
AREA=3.14*(100/2)^2
AREA=3.14*50^2
AREA=3.14*2,500
AREA=7,850 MI.^2 IS THE AREA.